Suppose a(x) = 3x - 7 and b(x) = 2 - x^3. Find a(b(3)) - b(a(3)).

Answer by robertb(5830)

b(3) = 2 - 3^3 = 2 - 27 = -25 ==> a(b(3)) = a(-25) = 3*-25 - 7 = -75 - 7 = -82
a(3) = 3*3 - 7 = 9 - 7 = 2 ==> b(a(3)) = b(2) = 2 - 2^3 = 2 - 8 = -6.
==> a(b(3)) - b(a(3)) = -82 - -6 = -76.